OpenStudy (anonymous):

Identify the 16th term of a geometric sequence where a1 = 4 and a8 = −8,748.

OpenStudy (anonymous):

in geometric sequence you multiply the previous number with a constant r right? so a1=4 a2=4*r a3=4*r^2 a4=4*r^3 ... a8=4*r^7 ... a16=4*r^15

OpenStudy (anonymous):

-8,748=4*r^7 find r and find 4*r^15

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

−172,186,884 −57,395,628 57,395,628 172,186,884

OpenStudy (anonymous):

it will be negative right?

OpenStudy (anonymous):

i need a calculator sec :D

OpenStudy (anonymous):

yes negative for sure

OpenStudy (anonymous):

which one?

OpenStudy (anonymous):

r is -3 right? find 4*(-3)^15 which is -57 million something

OpenStudy (anonymous):

A bacteria culture begins with 15 bacteria which double in amount at the end of every hour. How many bacteria exist in the culture at the end of the 12th hour?

OpenStudy (anonymous):

at first we got 15 lets call it b1, b1=15 b2=15 times 2 right? so b2=15*2 b3=15*2^2 so b12=15*2^11 right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

whats the final answer?

OpenStudy (anonymous):

@AllTehMaffs

OpenStudy (anonymous):

@AllTehMaffs

OpenStudy (anonymous):

The information included in this question is the intellectual property of Florida Virtual School. Posting questions from any FLVS course and receiving answers is considered cheating.